# Navigating the phase diagram of quantum many-body systems in phase space

### v1

@@ V Conclusion
We have analyzed the phase diagrams of spin$-(\frac{1}{2}-\frac{1}{2})$ and spin$-(\frac{1}{2}-1)$ Ising-Heisenberg chains through the lens of the Wigner function. Our results have illuminated the distinct roles played by the negative and positive parts of the Wigner function in revealing the complex phase structures and boundaries of these chains. The comparative analysis with entanglement concurrence has not only validated our findings but also highlighted the strengths and limitations of each approach in revealing the phase diagram of the systems.
The versatility of the equal angle slice approximation in the homogeneous spin$-(\frac{1}{2}-\frac{1}{2})$ chain, as opposed to the necessity of a full phase space integration for the inhomogeneous spin$-(\frac{1}{2}-1)$ chain, underlines the crucial influence of the quantum system homogeneity on phase space analysis. These insights pave the way for more nuanced approaches to studying quantum systems, particularly in understanding the interplay between system properties and the choice of phase space methods.
@@ Acknowledgements
Z.M. acknowledge support from the National Science Center (NCN), Poland, under Project No.${}2020/38/E/ST3/00269$. S.D. acknowledges support from the John Templeton Foundation under Grant No. 62422. B.G. acknowledge support from the National Science Center (NCN), Poland, under Project No.${}2022/47/B/ST6/02380$. This research is supported through computational resources of HPC-MARWAN provided by CNRST.
@@ Appendix A Spectrum of the ATIH chain
|$\ket{\phi_{1}}$ $=\frac{1}{\sqrt{1+e_{1}^{2}}}(e_{1}\ket{++}+\ket{--}),$ (21)
|$\ket{\phi_{2}}$ $=\frac{1}{\sqrt{2}}(\ket{+-}+\ket{-+}),$ (22)
|$\ket{\phi_{3}}$ $=\frac{1}{\sqrt{2}}(-\ket{+-}+\ket{-+}),$ (23)
|$\ket{\phi_{4}}$ $=\frac{1}{\sqrt{1+e_{2}^{2}}}(e_{2}\ket{++}+\ket{--}),$ (24)
@@
where the product is carried over all sites. The first element in the product corresponds to the Ising edge which can take only two possible values ±$\pm 1$, while the second element represents the Heisenberg edge. The Quantum Ferromagnetic phases are categorized into four types (QFO I-IV), each characterized by unique ground state vector products over all lattice sites, and they differ primarily in the orientation probabilities of spins in the Heisenberg edges, which are determined by the functions $e_{1}$ and $e_{2}$, Eq. (26). These functions highlight the quantum nature of the ferromagnetic state by indicating that up and down orientations have distinct probabilities. Meanwhile, the Frustrated Ferromagnetic states (FRU I-IV) demonstrate non-degenerate characteristics with spin frustration present, indicative of the competition between different interactions within the lattice.
@@
$\epsilon_{\pm}^{1}(\sigma_{i},\sigma_{i+1})$ =±$=\pm 2\alpha+\gamma+J_{z},$ (36)
$\epsilon_{\pm}^{2}(\sigma_{i},\sigma_{i+1})$ =±$=\pm\alpha+\gamma+\frac{1}{2}J_{+},$ (37)
$\epsilon_{\pm}^{3}(\sigma_{i},\sigma_{i+1})$ =±$=\pm\alpha+\gamma-\frac{1}{2}J_{+},$ (38)
@@
[…]
@@
[…]
where the product is carried over all the sites. The first element in the product corresponds to the Ising edge which can take only two possible values ±$\pm 1$, while the second element represents the Heisenberg edge.
@@ Appendix B The ATIH chain thermodynamics
Figure 5: Correlation function of the ATIH chain in the spin$-(\frac{1}{2}-\frac{1}{2})$ case
@@
where $J_{\nu},\nu=\{x,y,z\},$ denotes the coupling strengths between the Heisenberg nodes in the ``$\nu"$ direction. Finally, the Ising-Heisenberg spin correlation function
@@
The one-body and two-body correlation functions are given by Equations 65 and 66, while the three-body correlation function is written as $\langle\sigma_{i}\sigma_{i+1}\sigma_{j}\rangle=\langle\sigma_{i}\rangle^{3}+\langle\sigma_{i}\rangle(1-\langle\sigma_{i}\rangle^{2})(e^{(i-j)/\xi}+e^{-1/\xi}(1+e^{(j-i)/\xi}))$. $\xi$ is the correlation length of the standard Ising chain given by [92]
@@
Fig. (5) and (6) shows the correlation functions of the ATIH chain, Eq. (II), for the spin$-(\frac{1}{2}-\frac{1}{2})$ and spin$-(\frac{1}{2}-1)$ case, with respect to the set of parameters $(x,y)$ Eq. (2) and Eq. (3), respectively.
@@
Figure 6: Correlation function of the ATIH chain in the spin$-(\frac{1}{2}-1)$ case

### v2

@@ V Conclusion
We have analyzed the phase diagrams of spin(1212)-($\frac$$\frac{1}{2}$) and spin(121)-($\frac${1}{2}-1) Ising-Heisenberg chains through the lens of the Wigner function. Our results have illuminated the distinct roles played by the negative and positive parts of the Wigner function in revealing the complex phase structures and boundaries of these chains. The comparative analysis with entanglement concurrence has not only validated our findings but also highlighted the strengths and limitations of each approach in revealing the phase diagram of the systems.
The versatility of the equal angle slice approximation in the homogeneous spin(1212)-($\frac$$\frac{1}{2}$) chain, as opposed to the necessity of a full phase space integration for the inhomogeneous spin(121)-($\frac${1}{2}-1) chain, underlines the crucial influence of the quantum system homogeneity on phase space analysis. These insights pave the way for more nuanced approaches to studying quantum systems, particularly in understanding the interplay between system properties and the choice of phase space methods.
@@ Acknowledgements
Z.M. acknowledge support from the National Science Center (NCN), Poland, under Project No.2020/38/E/ST3/00269~2020/38/E/ST3/00269. S.D. acknowledges support from the John Templeton Foundation under Grant No. 62422. B.G. acknowledge support from the National Science Center (NCN), Poland, under Project No.2022/47/B/ST6/02380~2022/47/B/ST6/02380. This research is supported through computational resources of HPC-MARWAN provided by CNRST.
@@ Appendix A Spectrum of the ATIH chain
[…]
@@
where the product is carried over all sites. The first element in the product corresponds to the Ising edge which can take only two possible values $\pm 1$, while the second element represents the Heisenberg edge. The Quantum Ferromagnetic phases are categorized into four types (QFO I-IV), each characterized by unique ground state vector products over all lattice sites, and they differ primarily in the orientation probabilities of spins in the Heisenberg edges, which are determined by the functions $e_{1}$ and $e_{2}$, Eq. (26). These functions highlight the quantum nature of the ferromagnetic state by indicating that up and down orientations have distinct probabilities. Meanwhile, the Frustrated Ferromagnetic states (FRU I-IV) demonstrate non-degenerate characteristics with spin frustration present, indicative of the competition between different interactions within the lattice.
@@
$\epsilon_{\pm}^{1}(\sigma_{i},\sigma_{i+1})$ $=\pm 2\alpha+\gamma+J_{z},$ (36)
$\epsilon_{\pm}^{2}(\sigma_{i},\sigma_{i+1})$ $=\pm\alpha+\gamma+\frac{1}{2}J_{+},$ (37)
$\epsilon_{\pm}^{3}(\sigma_{i},\sigma_{i+1})$ $=\pm\alpha+\gamma-\frac{1}{2}J_{+},$ (38)
@@
$|\text{FM}\rangle$ $=\prod_{k=1}^{N}|+,\varphi_{+}^{1}\rangle_{k},$ (51)
$|\text{FRI}\rangle$ $=\prod_{k=1}^{N}|-,\varphi_{+}^{1}\rangle_{k},$ (52)
@@
$|\text{FRU}\rangle$ $=\prod_{k=1}^{N/2}|+,\varphi_{-}^{4},-\varphi_{-}^{4}\rangle_{k},$ (58)
where the product is carried over all the sites. The first element in the product corresponds to the Ising edge which can take only two possible values $\pm 1$, while the second element represents the Heisenberg edge.
@@ Appendix B The ATIH chain thermodynamics
Figure 5: Correlation function of the ATIH chain in the spin(1212)-($\frac$$\frac{1}{2}$) case
@@
where $J_{\nu},\nu=\{x,y,z\},$ denotes the coupling strengths between the Heisenberg nodes in the $\nu$" direction. Finally, the Ising-Heisenberg spin correlation function
@@
The one-body and two-body correlation functions are given by Equations˜65 and 66, while the three-body correlation function is written as $\langle\sigma_{i}\sigma_{i+1}\sigma_{j}\rangle=\langle\sigma_{i}\rangle^{3}+\langle\sigma_{i}\rangle(1-\langle\sigma_{i}\rangle^{2})(e^{(i-j)/\xi}+e^{-1/\xi}(1+e^{(j-i)/\xi}))$. $\xi$ is the correlation length of the standard Ising chain given by [92]
@@
Fig. (5) and (6) shows the correlation functions of the ATIH chain, Eq. (II), for the spin(1212)-($\frac$$\frac{1}{2}$) and spin(121)-($\frac${1}{2}-1) case, with respect to the set of parameters $(x,y)$ Eq. (2) and Eq. (3), respectively.
@@
Figure 6: Correlation function of the ATIH chain in the spin(121)-($\frac${1}{2}-1) case